OpenStudy (anonymous):

\[6\sqrt[4]{4} - 7\sqrt[3]{4} + 2\sqrt[3]{4} - 3\sqrt[4]{4}\] \[6\sqrt[4]{4} - 7\sqrt[3]{4} + 2\sqrt[3]{4} - 3\sqrt[4]{4}\] @Mathematics

OpenStudy (anonymous):

r u kidding me? just collect like terms

OpenStudy (anonymous):

you know you're awfully rude.

OpenStudy (anonymous):

you know your awfully not trying

OpenStudy (zarkon):

Lagrange must have gotten up on the wrong side of the multiplier

OpenStudy (anonymous):

this guy is really testing my method

OpenStudy (anonymous):

i have tried. i was in remedial math &failed regular math. if i knew how to do it i would try on my own. i don't need you coming on mine &giving your input if it's not anything nice. not everyone can be as smart as you.

OpenStudy (anonymous):

who said im smart, im in remdial math a well

OpenStudy (anonymous):

im in what they call in community college : math 98

OpenStudy (anonymous):

okay, i'm gonna be the bigger person &not argue with you. i'd appreciate it if you didn't comment on my questions unless it's something about the answer or it's something nice.

OpenStudy (anonymous):

okay, man stop posting questions without even having tried them

OpenStudy (anonymous):

Why are you being an retriceat to someone that needs help with math, then? You seem to be in very similar boats.

OpenStudy (anonymous):

first of all i'm a girl. it's clear as daylight i am. &you can't tell me i haven't tried them because i actually have. i ended up getting them wrong.

OpenStudy (zarkon):

\[6\sqrt[4]{4} - 7\sqrt[3]{4} + 2\sqrt[3]{4} - 3\sqrt[4]{4}\] \[=6\sqrt[4]{4} - 3\sqrt[4]{4}- 7\sqrt[3]{4} + 2\sqrt[3]{4}\] \[=(6-3)\sqrt[4]{4} +(- 7+2)\sqrt[3]{4} =\cdots\]

OpenStudy (anonymous):

thank you for the help zarkon.